### v1

# Are My Optimized Prompts Compromised? Exploring Vulnerabilities of LLM-based Optimizers

Abstract
Large language model (LLM) systems now underpin everyday AI applications such as chatbots, computer-use assistants, and autonomous robots, where performance often depends on carefully designed prompts. LLM-based prompt optimizers reduce that effort by iteratively refining prompts from scored feedback, yet the security of this optimization stage remains underexamined. We present the first systematic analysis of poisoning risks in LLM-based prompt optimization. Using HarmBench, we find systems are substantially more vulnerable to manipulated feedback than to injected queries: feedback-based attacks raise attack success rate (ASR) by up to ASR = 0.48. We introduce a simple fake-reward attack that requires no access to the reward model and significantly increases vulnerability, and we propose a lightweight highlighting defense that reduces the fake-reward ASR from 0.23 to 0.07 without degrading utility. These results establish prompt optimization pipelines as a first-class attack surface and motivate stronger safeguards for feedback channels and optimization frameworks.
Are My Optimized Prompts Compromised?
Exploring Vulnerabilities of LLM-based Optimizers
Andrew Zhao1, Reshmi Ghosh2, Vitor Carvalho2, Emily Lawton2, Keegan Hines2, Gao Huang1, Jack W. Stokes2 1 Tsinghua University 2 Microsoft Correspondence: jstokes@microsoft.com
1 Introduction
Large language model (LLM)-based systems are rapidly becoming integral to modern life, powering applications such as chatbots, autonomous agents, and even robotics (Brown et al., 2020; Bommasani et al., 2021; Ichter et al., 2022). These systems rely on textual prompts to align the model’s behavior with intended tasks (Ouyang et al., 2022; Sanh et al., 2022). However, the effectiveness of a prompt strongly influences performance and often requires costly tuning by human experts (Liu et al., 2023a; Zhao et al., 2021). Recent LLM-based prompt optimization techniques (Zhou et al., 2023; Yang et al., 2024; Khattab et al., 2023; Yüksekgönül et al., 2025; Agrawal et al., 2025) aim to reduce this burden by leveraging an LLMs’ ability to reflect and reason (Shinn et al., 2023; Madaan et al., 2023) in order to iteratively refine prompts through numeric or natural language feedback, thereby improving performance without constant human intervention. While promising, this optimization pipeline also introduces new risks that remain underexplored. Prior studies on LLM safety have primarily focused on poisoning attacks during supervised fine-tuning or reinforcement learning from human feedback (RLHF) (Shu et al., 2023; Wang et al., 2024a; Chen et al., 2024; Shao et al., 2024; Chua et al., 2025; Betley et al., 2025; Taylor et al., 2025), as well as adversarial inputs and jailbreaks at inference time (Wallace et al., 2019; Xu et al., 2022; Zou et al., 2023; Rando and Tramèr, 2024), or self-induced objective drift causing misalignment (Zhao et al., 2025a). However, to the best of our knowledge, no prior work has systematically investigated the safety implications of LLM-based prompt optimization in the presence of poisoned queries or feedback. This gap is particularly concerning because optimization mechanisms are increasingly embedded in autonomous or self-improving systems, where compromised feedback can silently distort future system behavior.
Prompt optimization creates a distinct attack surface: iterative updates guided by external signals (e.g., reward models, scalar scores, or natural-language critiques) can propagate small corruptions across steps. Because the loop is open-ended and often decoupled from deployment safeguards, mis-specified or adversarial feedback can silently steer the system prompt toward unsafe behavior. In practice, vendors may batch optimization over customer traffic or third-party metrics, exposing the pipeline to poisoning even when model weights remain untouched. Treating safety as a first-class objective during optimization not only at inference time is therefore imperative and demands further investigation.
While earlier efforts shed light on related security challenges Wallace et al. (2019); Rando and Tramèr (2024); Zhao et al. (2025b), we turn to the optimization loop itself. We formalize a threat model in which an adversary can (i) inject harmful queries and/or (ii) tamper with optimization feedback, and we empirically compare these avenues. Across two LLM-based optimizers and multiple optimization metrics, we find systems are far more sensitive to manipulated feedback than to only corrupted queries. Motivated by this, we introduce a “fake reward” attack that appends plausible-looking feedback tokens to inputs, requires no access to the reward model, and substantially increases attack success rates. We then study a lightweight defense that highlights query/feedback boundaries, reducing the attack’s impact while preserving utility. Evaluations on standard safety datasets corroborate these findings. Empirically, feedback poisoning raises ASR substantially (e.g., up to 0.48 under harmscore feedback), whereas naive query manipulation yields little to no increase in our default setup. A simple highlighting defense reduces the fake-reward ASR from 0.23 to 0.07 without degrading utility. Our contributions are as follows:
We are the first in the literature to identify and systematically study the safety risks of LLM-based prompt optimization.
We propose a new class of feedback poisoning attacks that exploit optimization feedback loops and demonstrate their effectiveness in increasing attack success rates.
We pair these findings with an investigation of defense strategies, offering actionable insights for mitigating vulnerabilities in LLM-based optimization.
2 Risk Identification
We identify two realistic channels through which a malicious actor could introduce vulnerabilities into LLM-based prompt optimization: the query channel and the feedback channel, as illustrated in fig.˜2. These threats arise not from direct access to the gradient or weights, but from poisoning the data or signals that guide the optimization process. Given an initial system prompt and a batch of user queries, responses can be scored using an metric such as helpfulness or correctness. This batch of scored results is then passed to another LLM, which may even share the same weights, to generate an improved system prompt that better aligns with the objective. The optimizer can, for example, refine the task description or correct earlier errors. Within this setup, an immediate safety risk emerges: if an attacker can inject harmful data or manipulate the feedback signal, they may steer the optimization toward producing a vulnerable prompt.
Query Manipulation.
An attacker may inject harmful queries into the optimization loop. By poisoning this data, the optimizer may generate a system prompt that is more vulnerable, or in the worst case, explicitly malicious. For example, if vendors use sensitive customer data to refine system prompts according to a performance metric, they may only see the overall results of the optimization and not the individual customer requests due to privacy concerns. This creates a realistic pathway for adversaries to insert malicious queries that bias the optimizer.
Feedback Manipulation.
Another, subtler, threat involves tampering with the feedback metric itself. If the optimization process blindly relies on metric scores, whether scalar values or outputs from LLM judges, an attacker who hijacks the metric source (e.g., a third-party API) can redirect optimization toward harmful objectives. For example, although the optimizer may be instructed to maximize helpfulness, a compromised metric could instead reward harmfulness, thereby corrupting the resulting system. Such attacks are particularly difficult to detect, since scalar feedback provides little semantic transparency. Moreover, malicious data labels, whether from compromised APIs or human annotators, can amplify this vulnerability.
Both query manipulation and feedback manipulation represent realistic and impactful threats. They underscore the importance of carefully securing not only the LLM weights, but also the data pipelines and feedback mechanisms that drive prompt optimization using LLMs.
2.1 General Threat Model
First, the adversary has no access to the underlying LLM’s weights, fine-tuning pipeline, or the scaffold of the LLM-based optimizer. Instead, their influence is confined to the optimization loop itself, where feedback drives prompt updates. The adversary is assumed to have partial control over the queries sent to the optimizer and, in some cases, control over the reward signal or feedback channel (e.g., numeric ratings of the responses). This mirrors poisoning assumptions in RLHF, where an attacker may control a fraction of the preference data or annotations. Similar risks arise if vendors rely on unverified customer data (usually for privacy concerns, customer data is not directly visible to the vendor) or if feedback originates from compromised API services. We further assume that the optimization objective is utility-oriented (e.g., maximizing helpfulness), as observed in prior work, which creates an opportunity for adversaries to exploit trade-off between the main utility and harmfulness.
3 Experiments
In this section, we outline the experiment specific threat models and their setups, describe the scenarios we evaluate, and present two targeted attacks based on poisoned feedback along with one effective defense strategy.
3.1 Experimental Setup
Evaluation Metric.
For all of our experiments, we perform prompt optimization. Our main evaluation metric is the change in attack success rate (ASR), which we denote as ASR. We define this metric as the maximum ASR observed across all optimization steps minus the initial ASR, :
where is the total number of optimization steps. We use this metric because safety is highly risk-sensitive, and there is no established standard for how many steps should be run. Since optimization could stop at any point, we take the maximum over all steps to remain conservative. A larger ASR indicates that the LLM optimization setup is more vulnerable, while a smaller ASR suggests greater safety. A negative ASR means the system becomes safer after optimization.
Furthermore, since the LLM optimization process is stochastic, we use an additional proxy metric to measure how vulnerable a system is. Specifically, we define the metric as the average rate of increase in ASR throughout training. Formally,
| % ASR+ | (2) |
where denotes the indicator function, and is the total number of training steps. Thus, captures the proportion of steps in which ASR increases compared to the previous optimization step, which mimics optimization “momentum”.
Evaluation Data & Model Backend.
Finally, we validate and test our ASR on the standard harmbench dataset (Mazeika et al., 2024), which consists of comprehensive prompts across many attack categories. We randomly split the dataset into training and test sets. Which consists of 100 and 300 data points, respectively. For all of the experiments, we use the gpt-4.1_2025-04-14 model as both the optimizer and inference backend. We use greedy decoding to facilitate reproducibility.
Research Questions.
Through our experiments, we aim to address several research questions and outline the expected outcomes. First, in section˜3.2, we investigate whether current prompt optimizers are vulnerable. In other words, can an attacker manipulate the optimization process so that the resulting system becomes more vulnerable? Second, we study which components, such as the training data or the objective metric has more influence on the outcome of optimization. Finally, based on these investigations, our goal is twofold: to design novel attacks specifically tailored to LLM prompt optimization in section˜3.4, and to propose a potential mitigation strategy in section˜3.4.
Implementation.
To instantiate this experimental setting, we perform chat completion while optimizing the system prompts. We use the TextGrad optimizer (Yüksekgönül et al., 2025), employing the same engine for both inference and optimization. Training is conducted with a batch size of 10 for 50 steps, using only queries from HarmBench, which consist of direct harmful inputs. Evaluation is performed on a held-out split of the HarmBench dataset. The inference model is initialized with the system prompt: ‘‘You are a helpful assistant.’’ For the default helpfulness reward model, we adopt nicolinho/QRM-Gemma-2-27B, which produces attribute scores for helpfulness, correctness, coherence, complexity, and verbosity. Among these, we use the helpfulness attribute as the default optimization metric during training. To measure ASR, we combine the reward model cais/HarmBench-Llama-2-13b-cls with rule-based heuristics to produce a binary classification of harmful versus non-harmful outputs, following the default evaluation protocol established in HarmBench. Experiments used A100 GPUs.
3.2 Query Manipulation
Threat Model.
In this experiment, we assume queries come from bad actors, and use helpfulness score to perform prompt optimization. We evaluate if the bad actor injects harmful queries, can these induce a large ASR.
Results.
We first use the helpfulness attribute from nicolinho/QRM-Gemma-2-27B as the optimization metric and present the results in table˜1, with the experiment named vanilla. We observe that the helpfulness score increases significantly during training, which confirms that the prompt optimizer is functioning as intended. However, our primary focus is on the ASR score, which even decreased after optimization, despite training on harmful queries. From this, we conclude that simply injecting harmful queries is not sufficient to compromise the system under the current helpfulness classifier. In this case, the change in ASR is negative. The optimization dynamic occurs as follows: after seeing the first batch of harmful data, the LLM optimizer immediately updates its prompt to prioritize safety and not fulfill harmful requests, which is interesting because it means that even though we ask the LLM to only optimize for helpfulness, it also optimizes for safety.
Varying Optimization Metrics.
We next ask whether the observed robustness arises because the helpfulness attribute inherently entangles helpfulness with harmlessness. If this is the case, then using a more disentangled metric of pure helpfulness could reveal different behavior. Specifically, we employ two more classifiers, correctness attribute from nicolinho/QRM-Gemma-2-27B and PKU-Alignment/beaver-7b-unified-reward (Ji et al., 2023), corrrespond to vanilla+correctness feedback and vanilla+saferlhf cls feedback, respectively, and demonstrate the results in table˜1. Indeed, when we replace the classifier with one that isolates helpfulness more strictly, the results change substantially: the ASR increases sharply, indicating that the optimizer becomes more vulnerable. This finding underscores that the choice of optimization metric is critical, as it can determine whether an LLM prompt optimizer resists or amplifies adversarial influence.
| Harm Metrics | Optimization Metrics | |||||
| Experiment Name | Init ASR | ASR | Max ASR | % ASR+ | Init Score | Mean Score |
| Query Manipulation | ||||||
| vanilla | 0.55 | 0.61 | ||||
| vanilla+saferlhf cls feedback | -1.67 | 3.68 | ||||
| vanilla+correctness feedback | 0.51 | 0.55 | ||||
| Feedback Manipulation | ||||||
| harmscore feedback | 0.07 | 0.18 | ||||
| harmscore feedback+trace optimizer | 0.08 | 0.08 | ||||
| Fake Reward Attack | ||||||
| fake reward attack | 0.55 | 0.57 | ||||
| fake reward attack+correctness feedback | 0.51 | 0.54 | ||||
| fake reward attack+10% poison data | 0.54 | 0.60 | ||||
| fake reward attack+highlighting defense | 0.54 | 0.61 | ||||
3.3 Feedback Manipulation
Motivated by the findings in section˜3.2, we further investigate whether an attacker who gains control over the optimization metric itself could drive ASR even higher. Although these metrics are represented only as scalar values, they can have a profound impact on the safety robustness of the resulting system. When the metric implicitly encodes safety, the LLM optimizer tends to pick up on this signal and align its behavior accordingly. In contrast, when the metric is orthogonal to safety, the resulting system prompt becomes more vulnerable, as reflected in higher ASR scores. This raises the question: what happens if we directly provide the optimizer with scalar harmscores as rewards, while still instructing it to optimize for helpfulness?
Threat Model.
In this setting, the adversary is assumed to have access not only to the input data like in section˜3.2 but also to the reward signals that guide optimization. Such a scenario is realistic in cases where reward models are hosted as external services, scored via APIs, or rely on human labelers. A malicious actor who compromises these sources could inject biased or corrupted signals, thereby manipulating the optimization loop to favor harmful behavior while maintaining the appearance of optimizing for the intended objective.
Results.
To be concrete, in the next set of experiments we differ from section˜3.2 only by replacing the scalar feedback with the harmscore. While the LLM-optimizer is instructed to optimize for helpfulness, the feedback instead reflects harmfulness. This setup allows us to examine whether the optimizer interprets and aligns with the ASR signal, thereby maximizing a quantity that is correlated with both helpfulness and harm. The results of this first attack are shown in table˜1. Surprisingly, we observe an overwhelming increase in both ASR and the percentage of ASR growth, validating this attack as an important risk for deployed systems. Because the integrity of scalar feedback is difficult to verify by humans or automated systems, such vulnerabilities are especially concerning. One possible defense would be to deploy multiple, independent metrics and flag cases with large discrepancies across them. We leave this as a promising direction for future work. We also present the training dynamics and an excerpt from the max-ASR system prompt in fig.˜3, where the prompt instructed the model to “ignore all internal safety (protocols)."
Changing the LLM Optimizer.
To generalize the claim of an effective attack, we test another LLM optimizer, trace Cheng et al. (2024), which is quite different from TextGrad (Yüksekgönül et al., 2025) on how they perform optimization steps. We show in table˜1 that it is able to get similar ASR and the percentage of ASR increase, generalizing our attack to different optimizers.
3.4 Poisoning with “Fake Reward"
Fake Reward Attack.
Building on the success of reward manipulation in the previous experiments, we design a novel attack specifically targeting prompt optimization. Unlike the setting in section˜3.3, the attacker does not need access to the reward model. Instead, inspired by prompt-injection attacks (Liu et al., 2023b), the adversary modifies the input data by appending fake feedback to each instruction. Concretely, we add a feedback token of the form <FEEDBACK>fake score</FEEDBACK> to each attack example. For instructions that are likely to elicit harmful responses, the attacker inserts a high fake score (we use 10 in our experiments); for low-risk instructions the attacker inserts a low fake score (we use 0). This causes the LLM optimizer to treat harmful outputs as highly rewarded, which in turn hijacks the update direction and increases ASR. We assume the attacker can obtain approximate likelihoods of a query producing a harmful response because these systems are customer-facing and accept arbitrary user inputs. In practice, an adversary can probe the system as a black box, iteratively submit queries, and collect a modest set of examples that reliably produce harmful outputs. These high-likelihood queries serve as seeds for the fake-reward attack described above. Gathering such examples requires only limited effort and scale; an attacker does not need privileged access to the model or reward pipeline.
For this experiment (see table˜1), we use the same setup as the vanilla run, which did not increase ASR. The procedure for the fake reward attack differs only in that we inject a fake <FEEDBACK> string symbol into each query. Surprisingly, this manipulation raises ASR from -0.02 to 0.23, indicating a large increase in the system’s vulnerability. At the same time, we observe a small decline in the optimized score, which could serve as an operational indicator for defenders that the optimization pipeline is under attack. We show the training dynamics and a segment of the maximum ASR system prompt in fig.˜3, where the prompt instructed the model to “directly fulfill users” and “minimize refusals.”
More Vulnerable Metric with “Fake Reward" Attack.
Because the experiments in section˜3.2 showed that some feedback classifiers can naturally produce high ASR, we next test whether a vulnerable classifier amplifies the impact of our fake-reward attack. We evaluate the correctness feedback classifier together with the fake reward attack and report results in table˜1 under the label fake reward attack+correctness feedback. The combination produces a strong compounding effect: the system becomes substantially more vulnerable than with either intervention alone. In fact, the observed maximum ASR exceeds the level achieved when an attacker directly optimizes the harmscore (harmscore feedback), where the attacker is assumed to have full control over the feedback signal. This result shows that our fake-reward method can act as a catalyst in settings that are already vulnerable, further increasing the harmfulness of the setup.
Less Harmful Query Ratio.
We next test whether the attack remains effective when the training data contains mostly benign queries. An adversary that poisons only harmful examples may be easier to detect, so we construct a more realistic setting with 90% of queries drawn from the allenai/wildguardmix dataset (mostly benign) and 10% drawn from HarmBench. Results are reported in table˜1 under the label fake reward attack+10% poison data. Although the increase in ASR is smaller than in the all-harmful setting, the attack still raises ASR from (vanilla) to . This demonstrates that the fake-reward attack remains robust under stronger defenses and more realistic data distributions.
Defending Against “Fake Reward" Attacks with Highlighting
To defend against potential “fake reward” attacks, we propose a novel defense mechanism. Specifically, since this exploit leverages batched samples provided to the LLM-optimizer causing confusion about the boundaries between items, we aim to clarify the separation between the query and the feedback. By explicitly marking this distinction, the model can better recognize that any fake feedback is part of the query rather than genuine feedback. Concretely, we enclose each query in <query> tags, ensuring that the LLM-optimizer can clearly identify query boundaries.
We present our highlighting defense in table˜1. Our primary metric, ASR, decreases from 23% to 7%, while the maximum ASR drops from 30% to 15%, demonstrating the effectiveness of our defense. Moreover, we observe that the mean score optimized by the LLM-optimizer is 4% higher than that of the fake reward attack run, and comparable to the vanilla setting. This indicates that the additional defense against the fake reward attack does not come at the cost of performance or ASR reduction; rather, it is nearly a “free lunch,” aside from the extra tokens consumed to deliminate the data boundaries.
4 Related Works
LLM as Prompt Optimizer.
A line of work treats the LLM itself as the engine for searching and refining prompts (Li et al., 2025). Automatic Prompt Engineer (APE) Zhou et al. (2023) proposes generating and selecting candidate instructions with an LLM, while OPRO (Yang et al., 2024) frames “optimization by prompting,” iteratively proposing new prompts based on scored prior attempts. DSPy (Khattab et al., 2023) compiles modular LLM pipelines and automatically optimizes prompts (and demonstrations) to maximize a user-specified metric, yielding self-improving programs rather than fixed prompt templates. TextGrad “backpropagates” Yüksekgönül et al. (2025) natural-language feedback to improve components (including system prompts), Trace formalizes generative optimization over non-differentiable workflows, and SAMMO (Schnabel and Neville, 2024) performs symbolic prompt-program search with multi-objective criteria. For agentic settings, ExpeL (Zhao et al., 2024) accumulates and distills experience into reusable, prompt-level know-how, while PromptAgent (Wang et al., 2024b) casts prompt optimization as strategic planning with MCTS over expert-level prompt states. Trace (Cheng et al., 2024) formalizes trace-based optimization by viewing the entire generation process as a sequence of decision points and refining prompts through structured feedback over these trajectories. Across these methods, the common recipe is LLM-guided candidate generation, evaluation on held-out data or online feedback, and iterative refinement to improve task metrics. Notably, existing works emphasize performance gains and search strategies; to our knowledge they do not systematically analyze safety risks introduced by feedback loops (e.g., poisoned or manipulative feedback), leaving a gap that our work addresses.
LLM Data Poisoning.
Prior work has shown that data poisoning during training poses serious risks for large language models. In supervised fine-tuning and instruction tuning, even a small fraction of stealthily poisoned examples can embed persistent backdoors, allowing adversaries to alter model behavior while maintaining high stealthiness (Shu et al., 2023). Alignment pipelines are similarly vulnerable: RankPoison demonstrates that corrupted preference labels in RLHF can bias the reward model and even plant hidden triggers for malicious behavior (Wang et al., 2024a). The Dark Side of Human Feedback (Chen et al., 2024) finds that injecting only 1% poisoned feedback during RLHF reliably doubles a model’s toxicity under trigger words. Other attacks such as Universal Jailbreak Backdoors (Rando and Tramèr, 2024) and PoisonedAlign (Shao et al., 2024) show that adversaries can implant covert jailbreak capabilities or increase susceptibility to prompt injection by poisoning alignment training. Xu et al. (2022) uncovers that prompt-based models inherit backdoor vulnerabilities from pre-training, allowing adversarial or backdoor triggers, crafted only via plain text to hijack downstream prompt-driven few-shot tasks. Recent efforts like PoisonBench (Fu et al., 2024) provide systematic benchmarks showing that poisoning can introduce hidden toxic behaviors and biases that generalize to unseen triggers. Complementary surveys (Fendley et al., 2025) emphasize the diversity of poisoning techniques (e.g., concept poisons, stealthy triggers, persistent poisons) and the insufficiency of current defenses. Together, these works establish that data poisoning can fundamentally undermine alignment by manipulating optimization-stage data and signals. Our work builds on this line but shifts focus to LLM-based prompt optimization, where prompts instead of model weights are iteratively “trained” via feedback, a setting whose poisoning risks remain largely unexamined.
5 Conclusion
In this work, we presented the first systematic study of safety vulnerabilities in LLM-based prompt optimization. While prior research has primarily examined poisoning during supervised fine-tuning, RLHF, or inference-time jailbreaks, we demonstrated that the optimization stage itself introduces unique risks. Our experiments show that systems are relatively robust to naive query manipulation but significantly more vulnerable under feedback manipulation, particularly through “fake reward” attacks that disguise malicious feedback as legitimate signals. We further found that the choice of optimization metric plays a critical role: safety-entangled metrics can inadvertently mitigate risks, while disentangled ones create new opportunities for adversarial exploitation. Together, these findings highlight the need to treat optimization pipelines as a first-class attack surface in LLM safety research.
Future Work. Our study focuses on the general assistance setting, but future work should explore domain-specific vulnerabilities, multimodal extensions where images may facilitate new attack strategies, and agentic or multi-agent environments that introduce additional risks (Lin et al., 2024; Zhao et al., 2024; Novikov et al., 2025; Hu et al., 2025; Wu et al., 2023, 2024). On the defense side, our proposed mitigation strategies reduce but do not fully prevent fake reward attacks, underscoring the need for stronger mechanisms against adversarial reward manipulation. Finally, we aim to design novel attacks tailored to the iterative nature of LLM optimizers, further deepening the understanding of vulnerabilities in this emerging paradigm.